Solve by the elimination method.
7r - 4s =24
4r + 7s =88
: Solve by the elimination method.
7r - 4s =24
4r + 7s =88

Answer by Alan3354(1187)

Solve by the elimination method.
7r - 4s =24 eqn 1
4r + 7s =88 eqn 2
------------
49r - 28s = 168 eqn 1 x 7
16r + 28s = 352 eqn 2 x 4
Add
65r = 520
r = 8
-----
Sub into either eqn
7*8 -4s = 24
56 - 4s = 24
-4s = -32
s = 8